# Image Inpainting based on Visual-Neural-Inspired Specific Object-of-Interest Imaging Technology

### v2

@@ 4.3 Preliminary Theoretical Verification
To isolate and validate the core contribution of our work—the Specific Object-of-Interest Imaging (SIOI) module—we first conducted a controlled experiment using a simplistic subsequent inpainting network (Stage II: SRM+GDRM). It is crucial to emphasize that this inpainting network is intentionally simplistic and is not the focus of our innovation. Its sole purpose is to serve as a stable and transparent baseline to demonstrate that the high-quality output is primarily attributable to the superior structural and semantic priors generated by the SIOI module (Stage I).
Unlike conventional image inpainting methods, our model eliminates the need for additional manually annotated corruption masks and is capable of automatically identifying and inpainting the corrupted regions of the target. We deliberately retained some black borders to demonstrate that the model does not solely rely on the color black to identify corrupted regions. As shown in Table 2 and Fig. 6, even with this simple network, the pipeline achieves exceptional performance on the Teapot dataset. This result provides foundational evidence that the SIOI module successfully accomplishes its design goal: to extract a clean, well-defined, and structurally sound representation of the target object from a complex and occluded scene. This output serves as an ideal input for anydownstream inpainting task, thereby validating the SIOI module’s efficacy as a standalone component.
@@ 4.4 Generalized empirical evaluation——Quantitative comparisons
Having established the standalone effectiveness of the SIOI module, we now evaluate its primary value proposition: its utility as a universal, plug-and-play front-end that can enhance diverse, off-the-shelf inpainting models. We integrated our SIOI module with a wide spectrum of state-of-the-art methods, spanning from classical algorithms to modern deep learning approaches. To ensure a fair comparison, all baseline methods were re-evaluated using a consistent object-level evaluation protocol, where their outputs were multiplied by corresponding object masks prior to quantitative analysis.
As summarized in Table 3, the SIOI module consistently enables significant performance improvements across all evaluation metrics. The key observation is not the absolute performance of any single model, but the transformative gain achieved by prepending SIOI. For instance, SIOI enables Criminisi to achieve a PSNR of 31.552 dB, a 7 dB improvement over the vanilla version. More notably, SIOI boosts LaMa to a staggering 45.222 dB PSNR and 0.0023 LPIPS, significantly surpassing all other methods. The improvements are especially pronounced on the Teapot dataset, where SIOI+LaMa achieves near-perfect reconstruction (SSIM: 0.9973, LPIPS: 0.0023).These results demonstrate that SIOI is not merely complementary but revolutionary, allowing both traditional and modern algorithms to transcend their inherent limitations by providing a purified, object-centric input. Regarding computational efficiency, SIOI acts as a generic adapter, effectively bridging the gap between holistic image processing and efficient object-centric reconstruction.
Across multiple datasets, models incorporating SIOI consistently achieve superior results in SSIM, PSNR, MAE, and LPIPS metrics compared to those without SIOI. Furthermore, models employing SIOI exhibit significantly enhanced perceptual quality. These results substantiate the effectiveness of SIOI in achieving high-fidelity reconstruction of objects with complex textures. Additionally, the object-centric imaging paradigm facilitates consistent separation of target objects across diverse backgrounds. Regarding computational efficiency, SIOI reduces redundant computations by concentrating processing resources on target regions, thereby accelerating inference while maintaining reconstruction quality.
@@
SIOI+Criminisi 0.9841 31.552 0.7396 0.0246
SIOI+PatchMatch 0.9752 25.912 1.7894 0.0408
SIOI+Edge-Connect 0.9795 32.378 0.8364 0.0299
SIOI+AOT-GAN 0.9248 36.940 1.0471 0.0094
SIOI+MISF 0.9739 33.487 1.0541 0.0674
SIOI+LaMa 0.9973 45.222 0.1119 0.0023
SIOI+RePaint 0.9712 28.599 1.6907 0.0645
@@
SIOI+Criminisi 0.9774 33.152 0.8312 0.0190
SIOI+PatchMatch 0.9721 29.327 1.3736 0.0325
SIOI+Edge-Connect 0.9636 30.693 1.2479 0.0448
SIOI+AOT-GAN 0.9341 34.691 1.4350 0.0243
SIOI+MISF 0.8883 29.070 3.4237 0.1848
SIOI+LaMa 0.9786 34.706 0.6523 0.0140
SIOI+RePaint 0.9102 28.956 3.2348 0.1489
@@
SIOI+Criminisi 0.9788 29.824 1.1000 0.0170
SIOI+PatchMatch 0.9734 26.327 1.7493 0.0306
SIOI+Edge-Connect 0.9676 27.527 1.6295 0.0382
SIOI+AOT-GAN 0.8741 30.861 2.3573 0.0350
SIOI+MISF 0.9325 27.270 3.0461 0.1215
SIOI+LaMa 0.9796 31.023 0.9491 0.0129
SIOI+RePaint 0.9465 27.109 2.8830 0.0992
@@
SIOI+Criminisi 0.9750 27.389 1.6940 0.0211
SIOI+PatchMatch 0.9675 24.691 2.4553 0.0365
SIOI+Edge-Connect 0.9632 26.089 2.2360 0.0387
SIOI+AOT-GAN 0.9070 28.025 3.2725 0.0527
SIOI+MISF 0.8960 24.048 6.2437 0.1556
SIOI+LaMa 0.9772 28.811 1.3896 0.0144
SIOI+RePaint 0.9130 24.309 5.9439 0.1365
@@ 4.5 Generalized empirical evaluation——Qualitative comparisons
Qualitative results further corroborate the role of SIOI as a powerful pre-processing stage. By providing a clean, background-suppressed, and structurally coherent representation of the target object, SIOI enables even conventional algorithms to produce more semantically consistent and visually plausible results. Fig. 7, Fig. 8, Fig. 9, Fig. 10, Fig. 11, Fig. 12, Fig. 13, Fig. 14, presents qualitative results that highlight characteristic behaviors and failure modes of the benchmark models. Criminisi [18] fundamentally lacks semantic understanding in its non-local patch matching, leading to severe structural discontinuities and semantically inconsistent textures in complex scenes. Similarly, PatchMatch [78] remains constrained by hand-crafted features and low-level metrics, failing to capture high-level semantics or global coherence, which results in visible patch artifacts—a byproduct of its localized/static patch-level processing that becomes evident under gradient-rich lighting and on non-rigid surfaces. AOT-GAN [79] exhibits jagged artifacts and discontinuous textures in complex regions, these failures stem from its use of dilated convolutions for feature aggregation. Edge-Connect [56], while effective in many cases, suffers from visible spatial misalignment due to its over-reliance on imperfect Canny edges, which do not correspond to perceptually relevant structures and disrupt spatial coherence in repetitive textures (e.g., grass). MISF [80] produces overly smooth, artificial-looking edges and illumination inconsistencies, resulting from its multi-scale regularization and strong smoothing priors as well as incompatible traditional and deep feature interactions. LaMa [81] over-smooths detailed regions and produces blurry or insufficiently sharp outputs despite its efficient Fourier-based design. RePaint [41] often yields overly smooth or blurry outputs with incongruous textures due to its slow. Without using our method, existing methods exhibit characteristic failures due to background interference and a lack of object-level understanding, resulting in color and texture details related to the background that are completely inconsistent with their original objects.
In contrast, the integration of the SIOI module consistently guides the inpainting focus, resulting in improved edge sharpness, texture consistency, and semantic coherence, regardless of the backend inpainting algorithm used. Experimental results across multiple object datasets show that our approach performs particularly well on rigid objects (e.g., teapots), while achieving relatively weaker results on non-rigid objects such as giraffes, zebras, and elephants. We attribute this difference to the higher morphological variability and structural complexity of non-rigid categories, which may require larger or more adaptive model parameters to adequately capture their semantic and textural characteristics. Preliminary observations suggest a potential linear relationship among semantic complexity, training data scale, and model parameter quantity—a phenomenon that merits further quantitative investigation in future work.
@@ 4.6 Ablation Studies
In fact, a detailed ablation experiment has already been conducted above, So this is just a simple supplement to the Preliminary Theoretical Verification experiment, also conducted on the Teapot dataset. This experiment is based on the preliminary theoretical verification experiment. To evaluate the effectiveness of our methodology, we conducted ablation studies under identical experimental conditions. Key components were systematically removed, and the resulting imaging outcomes were comparatively analyzed to elucidate performance differences. As shown in Fig. 15.
@@
Ablation I: In Ablation I, the state I network was removed while retaining the object-centric input–output pipeline. The corresponding results are presented in Tab. 4. The resulting model showed marginal superiority in certain quantitative metrics compared to the proposed method—SSIM (0.976 vs. 0.978), PSNR (33.95 dB vs. 33.86 dB), MAE (1.313 vs. 1.605), and LPIPS (0.017 vs. 0.018)—but these differences were statistically insignificant in terms of perceptual quality. However, subjective evaluation revealed compromised edge localization, manifested as jagged artifacts and global blurring. This degradation confirms the critical role of Stage I in achieving high structural precision. Furthermore, visual comparisons clearly demonstrated the superior inpainting quality achieved when Stage I was integrated. The marginal improvements observed in PSNR and MAE, despite perceptual and qualitative degradation, further highlight that such metrics can be misleading indicators of inpainting quality, as they from the loss function’s direct penalization of RGB deviations, do not consistently align with human perceptual judgment.
Ablation II: In Ablation II, the object-centric I/O was replaced with full-scene inputs. The model processed the entire image, including the extensive and complex background, which resulted in severe performance degradation. This represents the standard paradigm used by most existing methods, which our SIOI approach aims to supersede for object-centric tasks. Quantitative metrics showed substantial deterioration: SSIM: 0.964, PSNR: 29.43 dB, MAE: 2.085, LPIPS: 0.065. Visual inspection confirmed that this configuration yielded the poorest inpainting quality among all compared models. It exhibited pronounced detail loss, extensive artifacts throughout the restored regions, and markedly impaired image clarity. Furthermore, the model’s inability to effectively extract features, particularly from the edges of degraded regions, resulted in insufficient feature representation during inpainting. This shortfall directly led to blurry reconstruction and pervasive artifact generation. This ablation highlights the fundamental advantage of our object-first paradigm: it dramatically reduces the solution space by focusing computation exclusively on the target, thereby avoiding the pitfalls of whole-image processing.
In contrast to both ablated versions, the full model (with SIOI) achieved superior performance, delivering sharp edges, semantically coherent structures, and photorealistic textures. This provides compelling evidence that our two-stage pipeline, headed by the SIOI module, effectively addresses the bottlenecks of traditional inpainting by first resolving the ”where and what” of the target object before addressing the ”how” of its inpainting. The SIOI module acts as a powerful pre-processor that can be plugged into any inpainting network to elevate its performance on object-centric tasks.
Table 4: Comparison of ablation study results on the Teapot dataset. (mean $\pm$ std over five runs). ↑ means higher is better, and ↓ means lower is better.
@@ 4.7 Robustness Analysis
This experiment is based on the preliminary theoretical verification experiment, also conducted on the Teapot dataset. To evaluate the generalization ability of our method under challenging conditions, we conducted robustness experiments on the following scenarios:
@@
It should be noted that these challenging scenarios were not part of the training data. For qualitative assessment, two representative samples were selected per category. Quantitative evaluation was omitted due to the limited number of available samples for each specific condition, as illustrated in Fig. 16.
The qualitative results demonstrate the SIOI module’s notable robustness in several challenging conditions. It consistently maintained satisfactory target extraction performance under low illumination, small targets, salt-and-pepper noise, and motion blur. This resilience can be attributed to the module’s design principles: its spectral processing (FFC) is inherently robust to global illumination changes and motion artifacts, while its attentional mechanisms, inspired by V4 cortical processing, enable it to amplify and track small or salient targets despite noise. Performance degradation emerged in scenarios with three or more coexisting objects, as the competition for attentional resources increased. This observation directly mirrors the biological phenomenon of attentional limitation described in Desimone and Duncan’s biased competition model [66], validating the bio-inspired nature of our approach even in its failure modes.
The most substantial degradation occurred under high-power Gaussian noise. This is a recognized challenge for feature-based attentional systems both in biological and artificial networks. The SIOI module’s reliance on gradient and frequency information for target selection makes it particularly susceptible to this type of non-structured, high-entropy noise, which overwhelms the salient features the module is designed to detect. This result precisely highlights the critical role and inherent vulnerability of the front-end SIOI module: its output quality is the foundational prior for all subsequent processing. This underscores the importance of our modular design—the failure is contained and diagnosed at the SIOI stage, preventing a cascading collapse of the entire system and clearly pointing to the component that requires improvement.
These findings provide a clear roadmap for evolving the SIOI module. Future work will focus on enhancing its robustness against unstructured noise, potentially by incorporating neuromorphic noise-adaptation mechanisms inspired by the auditory brainstem or retinal adaptive gain control. This will further solidify the SIOI module’s role as a robust, plug-and-play pre-processor for object-centric vision tasks.
In summary, the robustness analysis confirms that the SIOI module provides a reliable and highly effective object-specific prior across a wide spectrum of challenging conditions. Its performance profile—strong in structured degradation but susceptible to unstructured noise—precisely characterizes its operational envelope and offers valuable insights for its application and further development as a general-purpose biological vision-inspired front end.
Figure 16: Illustration of robustness experiment.
@@ 5 Discussion
Image impainting is a job without a standard answer and often leads to controversy, as shown in Fig. 17. We don’t even know if the black area should be an buffalo head or a zebra head. The same model, when the same image is input multiple times, will yield different results. And the work proposed in this article can solve this problem. This work introduces Specific Object-of-Interest Imaging (SIOI)—a novel, bio-inspired paradigm that fundamentally rethinks the pipeline for object-centric visual computing tasks. Departing from conventional holistic processing, SIOI provides a plug-and-play front-end module that first isolates and enhances the target object, generating a refined structural and semantic prior that can be seamlessly integrated into any subsequent processing network. Our experimental validation demonstrates that incorporating the SIOI module consistently elevates the performance of diverse inpainting architectures. This is not a contest between inpainting models; rather, it is a demonstration that providing a high-quality, object-centric prior is a more impactful lever for improvement than designing yet another holistic inpainting network.
Figure 17: Misunderstand of image impainting.
@@
This approach aligns with the ventral pathway’s ”object-before-background” processing strategy, where attentional mechanisms in areas like V4 selectively amplify target features while suppressing irrelevant background clutter[83]. Our SIOI module computationally instantiates this principle, thereby mitigating the fundamental issue of information redundancy that plagues whole-image methods. Compared to GAN- or diffusion-based approaches that operate on the entire scene, our method avoids propagating errors from background regions by focusing computation on the target, akin to V4’s role in contour integration and feature binding.
@@
The robustness of the SIOI module under extreme conditions (e.g., low illumination, noise) further underscores its utility as a reliable pre-processor. Its behavior is consistent with the object-based attention mechanism in biological vision, which dynamically modulates early visual areas (V1-V4) to maintain target salience amidst distraction. It is important to note that the performance of any subsequent model is contingent on the quality of the prior provided by SIOI. This dependency mirrors the dorsal-ventral synergy in the brain, where spatial localization guides object recognition. This highlights that advancing object-level pre-processing is a critical frontier in itself.
@@
We emphasize that the inpainting network presented in Stage II serves primarily as a proof-of-concept vehicle to validate the efficacy of the SIOI prior. our architecture may not represent the ultimate performance ceiling, and that future researchers may develop superior networks. The paramount innovation of this work is not this specific network, but the SIOI paradigm itself. This paradigm shift—from direct holistic inpainting to object-first prior generation—is our central contribution. The generalization of the SIOI module is inherently tied to the semantic definition of a ”specific object.” This is not merely a limitation but a defining characteristic and a research opportunity. But we must frankly admit that the generalization of this technology is not strong due to its focus on specific objects. And each semantic object represents a type of dataset, which also requires retraining the object imaging network. For example, whether cat is an object, or is it divided into jungle cat felis chaus, sand cat felis Margarita, black foot cat felis nigripes, and wild cat felis silvestris(which is also more refined as a specific object) needs further research. The module’s performance can be scaled and specialized, offering a practical pathway for technology transfer, allowing the SIOI paradigm to be precisely tailored to application-specific needs.
@@
Future work will explore integrating prefrontal-cortex-like task-driven modulation to enable dynamic reconfiguration of the SIOI module for different object categories and end goals, advancing toward truly adaptive and biologically plausible computational vision systems.

### v3

@@ 4.3 Preliminary theoretical verification
To validate our core contribution—the Specific Object-of-Interest Imaging (SIOI) module—we conducted a controlled experiment using a simplified inpainting network (Stage II: SRM+GDRM). This network is intentionally simplified and not our innovation focus. Its purpose is to demonstrate that high-quality outputs primarily stem from SIOI’s structural and semantic priors (Stage I).
Unlike conventional methods, our model automatically identifies and inpaints corrupted regions without manual mask annotation. We retained black borders to show the model doesn’t rely solely on color cues for corruption detection. As shown in Table 2 and Fig. 6, this pipeline achieves exceptional performance on the Teapot dataset. These results confirm the SIOI module successfully extracts clean, structurally sound object representations from complex, occluded scenes. This validates SIOI’s efficacy as a standalone component for downstream inpainting tasks.
@@ 4.4 Generalized empirical evaluation——Quantitative comparisons
We next evaluate SIOI’s utility as a universal plug-and-play front-end for enhancing diverse inpainting models. We integrated SIOI with state-of-the-art methods spanning classical to modern deep learning approaches. For fair comparison, all methods were evaluated using a consistent object-level protocol, with outputs multiplied by corresponding object masks.
As shown in Table 3, SIOI consistently improves performance across all metrics. The key finding is the transformative gain from prepending SIOI, not any single model’s absolute performance. For example, SIOI improves Criminisi’s PSNR by 7 dB to 31.552 dB. Notably, SIOI boosts LaMa to 45.222 dB PSNR and 0.0023 LPIPS, surpassing all other methods. Improvements are most pronounced on the Teapot dataset, with SIOI+LaMa achieving near-perfect reconstruction (SSIM: 0.9973, LPIPS: 0.0023). These results show SIOI enables both traditional and modern algorithms to transcend their limitations through purified, object-centric inputs. SIOI also improves computational efficiency by bridging holistic image processing and efficient object-centric reconstruction.
Across datasets, models with SIOI consistently outperform those without it across all metrics. Models with SIOI also exhibit enhanced perceptual quality. These results confirm SIOI’s effectiveness for high-fidelity reconstruction of complex-textured objects. The object-centric paradigm enables consistent target separation across diverse backgrounds. Computationally, SIOI reduces redundant processing by focusing on target regions, accelerating inference while maintaining quality.
@@
SIOI+Criminisi 0.9841(↑4.5%) 31.5520(↑28.3%) 0.7396(↓79.2%) 0.0246(↓74.5%)
SIOI+PatchMatch 0.9752(↑5.9%) 25.9120(↑22.1%) 1.7894(↓67.2%) 0.0408(↓67.7%)
SIOI+Edge-Connect 0.9795(↑3.6%) 32.3780(↑29.0%) 0.8364(↓75.2%) 0.0299(↓67.1%)
SIOI+AOT-GAN 0.9248(↓2.5%) 36.9400(↑45.8%) 1.0471(↓67.8%) 0.0094(↓89.1%)
SIOI+MISF 0.9739(↑3.9%) 33.4870(↑34.6%) 1.0541(↓71.6%) 0.0674(↓42.8%)
SIOI+LaMa 0.9973(↑4.7%) 45.2220(↑77.1%) 0.1119(↓96.4%) 0.0023(↓97.2%)
SIOI+RePaint 0.9712(↑4.1%) 28.5990(↑19.0%) 1.6907(↓58.9%) 0.0645(↓43.5%)
@@
SIOI+Criminisi 0.9774(↑1.1%) 33.1520(↑1.6%) 0.8312(↓38.2%) 0.0190(↓27.5%)
SIOI+PatchMatch 0.9721(↑1.3%) 29.3270(↑2.4%) 1.3736(↓34.6%) 0.0325(↓26.3%)
SIOI+Edge-Connect 0.9636(↑0.8%) 30.6930(↓4.0%) 1.2479(↓28.5%) 0.0448(↑5.9%)
SIOI+AOT-GAN 0.9341(↑11.0%) 34.6910(↑25.5%) 1.4350(↓70.9%) 0.0243(↓60.0%)
SIOI+MISF 0.8883(↑4.9%) 29.0700(↑22.2%) 3.4237(↓35.3%) 0.1848(↓3.1%)
SIOI+LaMa 0.9786(↑6.0%) 34.7060(↑33.2%) 0.6523(↓78.0%) 0.0140(↓85.2%)
SIOI+RePaint 0.9102(↑5.5%) 28.9560(↑22.2%) 3.2348(↓37.5%) 0.1489(↓11.2%)
@@
SIOI+Criminisi 0.9788(↑0.7%) 29.8240(↓0.5%) 1.1000(↓26.9%) 0.0170(↓7.6%)
SIOI+PatchMatch 0.9734(↑0.8%) 26.3270(↓4.9%) 1.7493(↓24.0%) 0.0306(↓16.8%)
SIOI+Edge-Connect 0.9676(↑0.0%) 27.5270(↓8.0%) 1.6295(↓2.7%) 0.0382(↑36.4%)
SIOI+AOT-GAN 0.8741(↓4.0%) 30.8610(↑15.4%) 2.3573(↓43.2%) 0.0350(↓9.8%)
SIOI+MISF 0.9325(↑3.6%) 27.2700(↑16.0%) 3.0461(↓34.8%) 0.1215(↑0.6%)
SIOI+LaMa 0.9796(↑3.9%) 31.0230(↑20.6%) 0.9491(↓63.4%) 0.0129(↓80.7%)
SIOI+RePaint 0.9465(↑4.3%) 27.1090(↑15.0%) 2.8830(↓36.0%) 0.0992(↓9.5%)
@@
SIOI+Criminisi 0.9750(↑0.8%) 27.3890(↓0.8%) 1.6940(↓24.1%) 0.0211(↓10.6%)
SIOI+PatchMatch 0.9675(↑4.5%) 24.6910(↑9.1%) 2.4553(↓40.8%) 0.0365(↓57.4%)
SIOI+Edge-Connect 0.9632(↑0.5%) 26.0890(↓6.5%) 2.2360(↓28.0%) 0.0387(↑47.1%)
SIOI+AOT-GAN 0.9070(↑7.3%) 28.0250(↑25.1%) 3.2725(↓66.0%) 0.0527(↑8.4%)
SIOI+MISF 0.8960(↑4.4%) 24.0480(↑12.3%) 6.2437(↓24.4%) 0.1556(↓2.5%)
SIOI+LaMa 0.9772(↑4.7%) 28.8110(↑18.6%) 1.3896(↓62.9%) 0.0144(↓78.9%)
SIOI+RePaint 0.9130(↑5.2%) 24.3090(↑12.8%) 5.9439(↓25.9%) 0.1365(↓8.9%)
@@ 4.5 Generalized empirical evaluation——Qualitative comparisons
Qualitative results confirm SIOI’s effectiveness as a preprocessing stage. SIOI provides clean, background-suppressed object representations, enabling conventional algorithms to produce more consistent and plausible results. As shown in Figures 7-14, benchmark models exhibit characteristic failure modes: Criminisi [18] lacks semantic understanding, causing structural discontinuities and inconsistent textures; PatchMatch [80] relies on hand-crafted features, producing visible artifacts due to localized patch processing; AOT-GAN [81] shows jagged artifacts from dilated convolutions; Edge-Connect [44] suffers spatial misalignment from over-reliance on Canny edges; MISF [82] produces oversmoothed edges and illumination inconsistencies; LaMa [83] oversmooths details, producing blurry outputs; RePaint [41] yields blurry outputs with incongruous textures. Without SIOI, methods fail due to background interference and lack of object understanding.
With SIOI, all methods show improved edge sharpness, texture consistency, and semantic coherence. Our approach excels on rigid objects (e.g., teapots) but shows relatively weaker results on non-rigid objects (giraffes, zebras, elephants). This difference stems from higher morphological variability in non-rigid categories, requiring larger or more adaptive models. This suggests a relationship between semantic complexity, data scale, and model size, warranting future study.
@@ 4.6 Ablation Studies
In fact, a detailed ablation experiment has already been conducted above, So this is just a simple supplement to the Preliminary theoretical verification experiment, also conducted on the Teapot dataset. We conducted additional ablation studies on the Teapot dataset under identical conditions. Key components were systematically removed to analyze performance differences, as shown in Fig. 15.
@@
Ablation I: Removing Stage I while retaining object-centric I/O (Tab. 4) showed marginal metric improvements (SSIM: 0.976 vs. 0.978, PSNR: 33.95 dB vs. 33.86 dB, MAE: 1.313 vs. 1.605, LPIPS: 0.017 vs. 0.018), but these were statistically insignificant. Subjectively, edge localization degraded with jagged artifacts and blurring, confirming Stage I’s critical role for structural precision. The metric improvements, despite perceptual degradation, highlight that pixel-based metrics can mislead due to their focus on RGB deviations rather than human perception.
Ablation II: Replacing object-centric I/O with full-scene inputs (standard paradigm) caused severe performance degradation: SSIM: 0.964, PSNR: 29.43 dB, MAE: 2.085, LPIPS: 0.065. Visual inspection showed poorest quality with detail loss, artifacts, and impaired clarity. The model’s inability to extract features from degraded regions led to blurry reconstruction and artifacts. This highlights our object-first advantage: focusing computation on the target avoids whole-image pitfalls.
In contrast, the full model (with SIOI) achieved superior performance with sharp edges, semantic coherence, and photorealistic textures. This confirms that our two-stage pipeline effectively addresses traditional inpainting bottlenecks by first resolving target object localization and identity before inpainting. SIOI serves as a powerful pre-processor that enhances any inpainting network for object-centric tasks.
Table 4: Comparison of ablation study results on the Teapot dataset. (mean $\pm$ std over five runs). ↑ means higher is better, and ↓ means lower is better (teapot dataset).
@@ 4.7 Robustness Analysis
This study extends the preliminary theoretical verification experiment on the Teapot dataset. To assess generalization under challenging conditions, we tested robustness in scenarios:
@@
Note that these scenarios were absent from training data. For qualitative assessment, we selected two samples per category. Quantitative evaluation was omitted due to limited samples per condition, as shown in Fig. 16.
Qualitative results show SIOI’s robustness in several challenging conditions. It maintained satisfactory performance under low illumination, small targets, salt-and-pepper noise, and motion blur. This resilience stems from its design: spectral processing (FFC) resists global illumination changes and motion artifacts, while attentional mechanisms, inspired by V4, amplify and track small targets despite noise. Performance degraded with three or more objects due to attentional resource competition. This observation directly mirrors the biological phenomenon of attentional limitation described in Desimone and Duncan’s biased competition model [68].
Severe degradation occurred under high-power Gaussian noise. This challenge is known for feature-based attentional systems in biological and artificial networks. SIOI’s reliance on gradient and frequency information makes it susceptible to non-structured, high-entropy noise, overwhelming salient features. This highlights SIOI’s critical role and shortcoming: its output quality underpins all subsequent processing. Our modular design should contains failures at SIOI stage, preventing system collapse and pinpointing improvements. These findings guide SIOI’s evolution. Future work will enhance robustness against unstructured noise using neuromorphic adaptations from auditory brainstem or retinal gain control. This will solidify SIOI’s role as a plug-and-play pre-processor for object-centric vision.
In summary, SIOI provides reliable object-specific priors across challenging conditions. Its profile—strong against structured degradation but susceptible to unstructured noise—defines its operational envelope and guides application as a bio-inspired front end.
Figure 16: Illustration of robustness experiment (teapot dataset).
@@ 5 Discussion
Image inpainting lacks a definitive standard solution and frequently yields contentious results, as illustrated in Fig. 17. It is often ambiguous whether a missing region should be inpainted as one object or another (e.g., a buffalo or zebra head). Furthermore, the same model can produce varying outputs for an identical input across different runs. Our proposed framework addresses this challenge by introducing the Specific Object-of-Interest Imaging (SIOI) paradigm—a bio-inspired approach that reconceptualizes the processing pipeline for object-centric vision tasks. Departing from conventional holistic processing, SIOI provides a plug-and-play front-end module that first isolates and enhances the target object, generating a refined structural and semantic prior for any subsequent processing network. Our experiments demonstrate that incorporating SIOI consistently elevates the performance of diverse inpainting architectures. Thus, our contribution is not to propose a superior inpainting network, but to demonstrate that providing a high-quality, object-centric prior is a more impactful strategy than developing another holistic model.
Figure 17: Illustration of challenges in image inpainting.
@@
This approach mirrors the ventral visual pathway’s ”object-before-background” processing strategy, where attentional mechanisms (e.g., in area V4) enhance relevant features while suppressing background clutter [85]. Our SIOI module computationally implements this principle, mitigating the core issue of informational redundancy inherent in whole-image processing. Unlike GAN- or diffusion-based methods that process the entire scene, our approach avoids error propagation from background regions by concentrating computation on the target object, analogous to the role of V4 in contour integration and feature binding.
@@
The robustness of SIOI under challenging conditions (e.g., low illumination, noise) further highlights its utility as a reliable pre-processing module. Its behavior is consistent with the object-based attention mechanism in biological vision, which dynamically modulates early visual areas (V1-V4) to maintain target salience amidst distraction. Notably, the performance of any downstream model is contingent on the quality of the prior provided by SIOI. This dependency parallels the dorsal-ventral synergy in the brain, wherein spatial localization guides object recognition. This highlights that advancing object-level pre-processing is a critical frontier in itself.
@@
We emphasize that the inpainting network in Stage II serves primarily as a proof-of-concept to validate the efficacy of the SIOI prior. Consequently, our architecture does not represent the performance ceiling, and future work may develop superior networks. The paramount innovation of this work is not this specific network, but the SIOI paradigm itself. This paradigm shift—from direct holistic inpainting to object-first prior generation—is our central contribution. The generalization capability of the SIOI module is inherently linked to the semantic definition of a ”specific object.” This is not merely a limitation but a defining characteristic and a research opportunity. We acknowledge that the current generalization capability is limited due to its focus on specific object categories. Each semantically distinct object category effectively constitutes a separate dataset, necessitating retraining of the object imaging network. The granularity of this definition—for instance, whether ”cat” constitutes a single category or should be divided into subspecies (e.g., Felis chaus, Felis margarita)—requires further investigation. However, the module’s performance can be scaled and specialized, providing a practical pathway for technology transfer and enabling precise tailoring to application-specific needs.
@@
Future work will explore integrating prefrontal-cortex-like, task-driven modulation to enable dynamic reconfiguration of SIOI for different object categories and goals, advancing toward adaptive and biologically plausible vision systems.